OpenStudy (anonymous):

only 2 problems left...6a ------- how do you do this? 5a^3b^3

OpenStudy (anonymous):

\[\frac{6}{5a^2b^3}\] using the same techniques from the previous problem i showed you

OpenStudy (anonymous):

Could you please be more specific with spacing and parentheses? There is a big difference between 5a^3, (5a)^3

OpenStudy (anonymous):

tyler, so would you minus 1 from 3 to get 2 on top?

OpenStudy (anonymous):

You would get \[\frac{6a^{-2}}{5b^3}\]

OpenStudy (anonymous):

You always move negative exponents to the bottom therefore you get \[\frac{6}{5a^2b^3}\]

OpenStudy (anonymous):

Well not always but its a good idea